Circular No. 6642

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1997 J1 Jean Mueller reports her discovery of a comet on a 40-min IIIa-J ex-

posure by J. D. Mendenhall and herself with the 1.2-m Oschin Schmidt telescope in the course of the second Palomar Sky Survey. The object’s trail is condensed, with a diffuse coma and a hint of a tail to the southeast. The discovery was confirmed on another 40-min exposure two nights later. The following positions of the trail ends were measured by Mueller and reduced by G. V. Williams:

1997 UT  α_{2000}  δ_{2000}  m_{1}
May 3.23542  11^{h}39^{m}24.^{s}56  +75^{o}42′45′′.5  14
3.26319  11 39 08.45  +75 42 19.7
5.27153  11 18 18.25  +75 01 22.0
5.29931  11 18 04.26  +75 00 49.8

SUPERNOVA 1997br IN ESO 576-40 W.-d. Li, Beijing Astronomical Observatory (BAO), reports for the

BAO Supernova Survey: “SN 1997br in ESO 576-40 (also known as ESO 576-G40) has been followed photometrically at the BAO with the 0.6-m reflector and spectroscopically with the 2.16-m telescope, confirming that SN 1997br is a peculiar (unusally bright) type-Ia object. Our observations show that the supernova was on rise from Apr. 10 to 24, reaching its maximum optical brightness at magnitude R ∼ 13.7 and declining thereafter; the slow brightening of SN 1997br (only 0.6 mag in 13 days) also suggests a photometric evolution similar to that of SN 1991T. The spectroscopic observations obtained on Apr. 11, 12, 13, 14, 16, and 28 also show an evolution similar to that of SN 1991T. The spectra of SN 1997br taken before its optical maximum show no conspicuous Si absorption, with Fe lines dominating. The data taken on Apr. 28 (∼ 4 days past maximum), however, show a spectrum typical of normal type-Ia supernovae; conspicuous and relatively strong Si absorption was seen at 628 nm, and the spectrum is dominated by lines of S, Si, Ca, and Mg. The expansion velocity of the envelope is estimated to be ∼ 8300 km/s, as derived from the Si absorption line at 628 nm.”

1997 May 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green